OpenStudy (anonymous):

Assuming that the density of vinegar is 1.005 g/mL, calculate the molarity of acetic acid in vinegar from your average value for the mass percentage of acetic acid in vinegar. Density of vinegar= 1.005 g/mL Average mass % acetic acid= 5.2%

OpenStudy (anonymous):

Calculate mass of Acetic Acid 1.005*5.2%=.052g Acetic Acid Calculate mol of Acetic acid from molar mass .052g/60g/mol=8.6*10^-4 mol Calculate molarity mol/L 8.6*10^-4mol/.001L= .86M